OpenStudy (anonymous):

Let A = {circle, square, triangle} and let B = {3, 4}. Find A X B. {circle, square, triangle, 3, 4} { } {(circle, 3), (circle, 4), (square, 3), (square, 4), (triangle, 3), (triangle, 4)} {3, 4}

OpenStudy (anonymous):

need help n explain to me plzz thxs very muuch

OpenStudy (anonymous):

the third one

OpenStudy (anonymous):

It's easy to visualize with the combination of pizza and toppings choice :)

OpenStudy (cwrw238):

every element in A is paired with those in B

OpenStudy (cwrw238):

good answer chlorophyll

OpenStudy (anonymous):

The "\(\times\)" Denotes the Cartesian product. So, if we have two sets \(A,B\) we take each element of \(A\) and "concatenate it" with each element of \(B\). @Chlorophyll 's way is the perfect way to start!

OpenStudy (anonymous):

or say in mathematical way branch tree :)

OpenStudy (anonymous):

soo it would be third one right??

OpenStudy (anonymous):

Or do tabulate with: row is shape and col is number, knowing the total result 3* 2 = 6 elements

OpenStudy (anonymous):

@neel001 Yep.

OpenStudy (anonymous):

(Thank you)^n, everybody is so generous :)

OpenStudy (anonymous):

yhxs very much and one more question

OpenStudy (anonymous):

If g(x) = x2 − 3, find g(4). 13 5 16 −3

OpenStudy (anonymous):

Open the new post for new question as the rule in this site, will you :)

OpenStudy (anonymous):

okay but will u help me